Solve
(D^2+5DD'+5D'^2)z= x.sin(3x-2y)
1
Expert's answer
2020-11-09T20:17:30-0500

Given "(D^2+5DD'+5D'^2)z= x.sin(3x-2y)"

The auxiliary equation is

"m^2+5m+5=0 \implies \frac{-b \pm \sqrt{b^2-4ac}}{2a} =\frac{-5 \pm \sqrt{5^2-4(1)(5)}}{2(1)} \\
i.e.\ m=-5,0\\"

whence complementary function (C.F.) =

"\phi_1(y-5x)+\phi_2(y)\\"

"P.I. = \frac{1}{D^2+5DD'+5D'^2}x \cdot sin(3x-2y)=
\begin{cases} 
 D^2=-9 \\
 DD'=6 \\
 D'=-4 
 \end{cases}"


(i.e., on replacing "D^2" by "a^2,\ DD'" by "-ab, D'^2=-b^2" provided "\u03c6(\u2013a2, \u2013ab, \u2013b2) \u2260 0" )

Clearly it is a case of failure as "\u03c6(\u2013a2, \u2013ab, \u2013b2) = 0."


Therefore, factorize "\phi (D,D')" and apply these factors turn by turn

"P.I.=\frac{1}{D(D+5D')}x\ sin (3x-2y)\\
=\frac{1}{D} \intop_{D+5D'}x\ sin (3x-10x-2c_1)\ \therefore\ y-5x=c_1\\
=\frac{1}{D} \intop x\ sin (-7x-2c_1)\delta x =\frac{1}{D} \intop \frac{- x\ cos (-7x-2c_1)}{-7}+\frac{1}{-7}\frac{sin (-7x-2c_1)}{-7}\\
=\frac{1}{D}\frac{x}{7}cos(3x-2y)+\frac{1}{49}sin (3x-2y)\\
= \intop_D(\frac{x}{7}cos(3x-2y)+\frac{1}{49}sin(3x-2y))\delta x\\

= \intop_D(\frac{x}{7}cos(3x-2c_2)+\frac{1}{49}sin(3x-2c_2))\delta x\ \therefore y=c_2\\"


"= \intop(\frac{x}{7}cos(3x-2c_2)\delta x+\intop\frac{1}{49}sin(3x-2c_2))\delta x\"

Letting "a= \intop(\frac{x}{7}cos(3x-2c_2)\delta x" and "b=\intop\frac{1}{49}sin(3x-2c_2))\delta x\"

"a= \intop(\frac{x}{7}cos(3x-2c_2)\delta x =\frac{x\ sin(3x-2c_2)}{21}+\frac13 \frac{cos(3x-2c_2)}{3}\\
a\implies \frac{x}{21}sin(3x-2c_2) +\frac{1}{9}cos(3x-2c_2)\\

b= \intop(\frac{1}{49}sin(3x-2c_2)\delta x =\frac{-cos(3x-2c_2)}{49}+\frac13 \frac{sin(3x-2c_2)}{49}\\
b\implies -\frac{1}{49}cos(3x-2c_2) +\frac{1}{147}sin(3x-2c_2)\\

\therefore a+b \implies \frac{x}{21}sin(3x-2c_2) +\frac{1}{9}cos(3x-2c_2)-\frac{1}{49}cos(3x-2c_2) +\frac{1}{147}sin(3x-2c_2)\\ 

\implies \frac{7x}{147}sin(3x-2c_2)+\frac{1}{147}sin(3x-2c_2) +\frac{40}{441}cos(3x-2c_2)\\ 
P.I. =\frac{7x}{147}sin(3x-2y)+\frac{1}{147}sin(3x-2y) +\frac{40}{441}cos(3x-2y)\\"

"\therefore" The complete Integral "z=C.I.+P.I."


Answer

"=\phi_1(y-5x)+\phi_2(y)+\frac{7x}{147}sin(3x-2y)+\frac{1}{147}sin(3x-2y) +\frac{40}{441}cos(3x-2y)\\"